Circular No. 7363

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SAGITTARII 2000 R. M. Wagner and R. Bertram, Ohio State University; and S. G.

Starrfield, Arizona State University, report: “On Feb. 11.55 UT, we obtained a spectrum (range 395–600 nm, resolution 1.3 nm) with the 2.4-m Hiltner Telescope of the MDM Observatory on Kitt Peak of the possible nova reported on IAUC 7362. The spectrum exhibits broad emission lines from the Balmer series of hydrogen and Fe II superposed on a red continuum. The Hβ and Hγ emission lines exhibit P-Cyg profiles. The instrument-corrected FWHM of Hβ emission is 1530 km/s, implying an expansion velocity of 765 km/s. The equivalent width of Hβ emission is 16 nm. The spectrum is typical of classical novae in the early decline phase. Additional observations are encouraged.”

XTE J1543–568 F. E. Marshall, Goddard Space Flight Center (GSFC); T. Takeshima,

GSFC and Universities Space Research Association; and J. in ’t Zand, Space Research Organization Netherlands, report the discovery of a transient, pulsating x-ray source designated XTE J1543-568: “The source was first seen in a slew of the Proportional Counter Array on the Rossi X-ray Timing Explorer at Feb. 6.84 UT with a 2–10-keV flux of 2 × 10^{−10} erg cm^{−2} s^{−1}. A subsequent observation with the PCA on Feb. 10.69 found a similar intensity and located the source in an error box that corresponds to galactic coordinates l = 324^{o}.88 ± 0^{o}.2, b = −1^{o}.45 ± 0^{o}.1, centered on α = 15^{h}43^{m}33^{s}, δ = −56^{o}48′.0 (equinox 2000.0). The flux is strongly modulated with a period of 27.12 ± 0.02 s. Observations at other wavelengths are encouraged.”

SAX J1808.4–3658 S. Wachter and D. W. Hoard, Cerro Tololo Interamerican Observa-

tory (CTIO), report: “We observed the optical counterpart (Giles et al. 1999, MNRAS 304, 47) of the x-ray transient and millisecond pulsar SAX J1808.4–3658 that was recently reported to exhibit renewed x-ray activity (IAUC 7358). On Feb. 9.38 UT, using the CTIO 0.9-m telescope (+ 2K CCD), we measured I = 18.2 ± 0.1 relative to the local standards given by Giles et al. (ibid.). This indicates that the counterpart has not reached the brightness levels (I = 15.6, IAUC 6885) observed during the previous outburst in 1998 (IAUC 6876).”

2000 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green